OpenStudy (anonymous):

Metallothermic production of zirconium (Zr) would involve the reaction of sodium (Na) with zirconium tetrachloride (ZrCl4) to produce zirconium and NaCl. (a) Write the balanced chemical equation using lowest whole number coefficients. (b) Calculate the amount of zirconium produced (in kg) if a reactor were charged with 194.0 kg ZrCl 4 and 19.0 kg Na .

OpenStudy (matt101):

Have you figured out the chemical equation?

OpenStudy (academicgurusinc):

For a) \[ZrCl4 + 4Na \rightarrow Zr + 4NaCl\] For (b) According to balanced chemical equation, one mole of ZrCl4 (0.233 Kg of ZrCl4) produces 1 mole Zirconium metal (0.091 Kg). We are provided with 194 Kg of ZrCl4 therefore, The amount of Zirconium produced will be = (194 * 0.091)/(0.233) = 75.77 kg of Zr metal approximately. In the same way one mole of 4 moles of Na metal (0.092Kg) will produce one mole of Zr metal (0.091Kg). Therefore, The amount of Zr metal produced by 19Kg of Na metal will be = (19*0.091)/(0.092) = 18.79Kg approximately. In this way Na is the limiting reactant and thus the theoratical yield will be dependent upon it i.e. 18.79Kg. I hope that helps. If you found this helpful, I encourage you to subscribe to our youtube channel ( https://www.youtube.com/channel/UCYiI7SmkU4_vhdSzKBWsifg), to stay current with all of our new videos. Regards, Academic Gurus Inc. Twitter (Academic_Gurus) Facebook (AcademicGurusInc) Youtube (Academic Gurus Inc)